OpenStudy (anonymous):

The equation of the half-value layer is d(1/2) = ln2/attenuation coefficient. How do i calculate the attenuation coefficient?

OpenStudy (shane_b):

Assuming this is your equation:\[\frac{d}{2}=\frac{ln(2)}{attenuation~coef}\]Then you just need to do a little bit of algebra to get:\[attenuation~coef=\frac{ln(4)}{d}\]